# A Theoretical Framework for Grokking: Interpolation followed by Riemannian Norm Minimisation

### v1

@@ B.4 Bounding the trajectories
If Assumptions 1 and 2 hold, there exists a compact $K$ of $\mathbb{R}^{d}$ such that for any $\lambda>0$ and $t\geq 0$, $w^{\lambda}(t)\in K$.
In particular, $\lim_{t\to\infty}w^{\lambda}(t)$ exists for any $\lambda>0$.
@@
Similarly to the proof of Lemma 1, we can consider a neighbourhood $U$ of $\mathcal{M}$ where the PL inequality holds:
@@
1) For some fixed $\varepsilon>0$, Lemma 2 then implies there is a $\lambda^{\star}>0$ and times $t(\lambda)$ such that for any $\lambda\in(0,\lambda^{\star})$ both hold
@@
Moreover, the proof of Lemma 2 also implies that there is some compact $K_{1}$ of $\mathbb{R}^{d}$ such that for any $\lambda<\lambda^{\star}$ and $t\leq\frac{t(\lambda)}{\lambda}$, $w^{\lambda}(t)\in K_{1}$.
@@
$\leq-\|\nabla F(w^{\lambda}(t))\|_{2}^{2}+\lambda\|w^{\lambda}(t)\|\|\nabla F(w^{\lambda}(t))\|$
@@
Define $t^{\prime}=\min(\frac{t(\lambda)}{\lambda}+\frac{-\ln(\lambda)}{\eta},T_{\lambda})$. Using Equation 18 for any $t\in(\frac{t(\lambda)}{\lambda},t^{\prime}]$:
@@
$\left\|w^{\lambda}(t)-w^{\lambda}(\frac{t(\lambda)}{\lambda})\right\|$ $\leq\int_{\frac{t(\lambda)}{\lambda}}^{t^{\prime}}\|\nabla F_{\lambda}(w^{\lambda}(s))\|\mathrm{d}s$
$\leq\beta\int_{\frac{t(\lambda)}{\lambda}}^{t^{\prime}}\sqrt{F(w^{\lambda}(s))-F^{\star}}\mathrm{d}s.$
From there, Equation 20 yields for any $t\in(\frac{t(\lambda)}{\lambda},t^{\prime}]$:
$\left\|w^{\lambda}(t)-w^{\lambda}(\frac{t(\lambda)}{\lambda})\right\|$ $\leq\beta\int_{\frac{t(\lambda)}{\lambda}}^{t^{\prime}}\sqrt{\varepsilon}e^{-\frac{\eta}{2}s}\mathrm{d}s+\beta(t^{\prime}-\frac{t(\lambda)}{\lambda})\sqrt{\lambda(\frac{R_{1}}{\eta}+\frac{1}{2}R^{2})}$
$\leq\beta\sqrt{\varepsilon}\frac{2}{\eta}-C\sqrt{\lambda}\ln(\lambda),$
@@
3) Since $t^{\prime}<T_{\lambda}$, $t^{\prime}=\frac{t(\lambda)}{\lambda}+\frac{-\ln(\lambda)}{\eta}$ by definition and
@@
Now define $\tilde{T}_{\lambda}=\inf\left\{t\geq\frac{t(\lambda)}{\lambda}\mid u^{\lambda}(t)\not\in U\right\}$. Since $\mathcal{M}$ minimizes $F$ on $U$, Equation 21 implies by continuity that for any $t\in[t^{\prime},\tilde{T}_{\lambda}]$:
@@
In particular, for $K_{2}=B(0,\frac{2R_{1}}{\eta}+R^{2}+3\varepsilon)$ and any $t\in[t^{\prime},\tilde{T}_{\lambda}]$, $w^{\lambda}(t)\in K_{2}$. By continuity and compactness, $\inf_{w\in(\partial U)\cap K_{2}}F(w)>F^{\star}$ thanks to Lemma 5. In consequence, we can choose $\lambda$ small enough so that Equation 21 implies that for any $t\in[t^{\prime},\tilde{T}_{\lambda}]$,
@@
[…]
To summarize, we have showed that there exists a small enough $\lambda^{\star}$, such that for any $\lambda<\lambda^{\star}$
@@
As a consequence, the definability assumption of $F_{\lambda}$ along with the boundedness implies that $\lim_{t\to\infty}w^{\lambda}(t)$ exists thanks to Kurdyka [1998, Theorem 2].
B.5 Proof of Lemma 3
@@
We consider the neighbourhood $U$ defined as in Lemma 7. By definition, $F$ is constant on the manifold $\mathcal{M}$ and denote its value $F^{\star}$, i.e., $F^{\star}=\sup_{w\in\mathcal{M}}F(w)=\inf_{w\in\mathcal{M}}F(w)$.
@@
Thanks to Lemma 5, for any $w\in U\setminus\mathcal{M}$, $F(w)>F^{\star}$. This implies the third condition,
@@
The arguments of Lemma 7 extend to any family of solutions $(u^{\lambda})_{\lambda>0}$ satisfying the assumptions of Lemma 3. In consequence, we can consider a compact $K$ of $\mathbb{R}^{d}$ such that for any $t\geq 0$ and $\lambda>0$, $u^{\lambda}(t)\in K$. From there, note again that $\inf_{w\in(\partial U)\cap K}F(w)>F^{\star}$.
@@
By monotonicity of the objective over time, $F(u^{\lambda}(t))+\lambda\|u^{\lambda}(t)\|^{2}_{2}<\min(F^{\star}+\varepsilon,\inf_{w\in(\partial U)\cap K}F(w))$ for any $t\geq 0$. Since $u^{\lambda}(t)$ is continuous, it implies that $u^{\lambda}(t)\in\overset{\circ}{U_{\varepsilon}}$ for any $t\geq 0$, which concludes the proof of Lemma 3. ∎
@@ B.6 Strict Minimality
Under the same assumptions than Proposition 4 with $w^{\star}=\lim_{t\to\infty}\tilde{w}^{\circ}(t)$, there exists a $\delta^{\star}>0$ such that for any $\delta\in(0,\delta^{\star})$, there exists $\varepsilon>0$ and $\lambda^{\star}>0$ such that for any $\lambda\in(0,\lambda^{\star})$
@@
Let $\delta^{\star}>0$ be such that for any $w\in\mathcal{M}\cap B(w^{\star},2\delta^{\star})$, $w\neq w^{\star}\implies\|w\|^{2}>\|w^{\star}\|^{2}$ and such that Equation 14 holds. Now let $\delta\in(0,\delta^{\star})$. We now fix $\lambda^{\star}>0$ arbitrarily small, choose $\lambda\in(0,\lambda^{\star})$ and define
@@ Appendix C Applications
In this section, we provide additional details to the examples discussed in Section 5, and specify how our theoretical results can be applied in various settings.
@@
Denote the solution of minimal $\ell_{2}$ norm with $w^{\star}=X^{+}y$, where $X^{+}$ is the Moore-Penrose pseudoinverse of $X$. The problem is convex and the critical set of $F$ is the affine subspace $\mathcal{M}=w^{\star}+\mathrm{Ker}(X)$, which is a manifold: Assumption 2 is satisfied.
@@
DLNs serve as a toy example to understand the influence of the architecture on the training dynamics of neural networks [Pesme, 2024]. The corresponding optimization problem writes
@@
$u\odot v\odot\left[X^{\top}(X(u^{2}-v^{2})-y)\right]=0,\$ (25)
@@
The set $\mathcal{M}^{*}=\nabla F^{-1}(0)\cap(\mathbb{R}^{*})^{2d}$ is a smooth manifold of dimension $2d-n$.
@@
Let $(\bar{u},\bar{v})\in\mathcal{M}^{*}$. Denote $W$ a neighborhood of $(\bar{u},\bar{v})$ such that $U\subset(\mathbb{R}^{*})^{2d}$. The function $H:\mathbb{R}^{2d}\rightarrow R^{d}$ with $H(u,v)=X^{\top}(X(u\odot v)-y)$ is a local defining function for $\mathcal{M}^{*}$, in the sense that for $(u,v)\in W$, we have $(u,v)\in\mathcal{M}^{*}\iff H(u,v)=0$.
@@
It is clear that, since all coordinates of $(\bar{u},\bar{v})$ are nonzero, the map $(\Delta u,\Delta v)\mapsto\bar{u}\odot\Delta u-\bar{v}\odot\Delta v$ is a surjection on $\mathbb{R}^{d}$, and therefore $\text{rank}(DH(\bar{u},\bar{v}))=\text{rank}(X^{\top}X)=n$. This proves that $\mathcal{M}^{*}$ is a manifold of dimension $2d-n$ [Boumal, 2023, §3.2]. ∎
Because of the singular points in $\mathcal{M}$, the function $F$ does not satisfy Assumption 2 globally. However, our results can still be applied locally: see the paragraph below for details.
@@
A typical example is symmetric matrix completion, where the goal is to recover an unknown matrix $M^{*}\in\mathbb{R}^{n\times n}$ from a subset of observed entries with coefficients in $\Omega\in\{1\dots n\}^{2}$: the objective function writes $F(W)=\sum_{(i,j)\in\Omega}\left((WW^{\top})_{ij}-M_{ij}^{*}\right)^{2}$. Note that the asymmetric case presented in Section 5, Equation 8, can also be written as a symmetric matrix completion problem, by setting
@@
Let $\overline{W}\in\mathcal{M}^{*}$, and let $\mathcal{U}$ a neighborhood of $\overline{W}$ such that $\mathcal{U}\subset\mathbb{R}^{n\times n}_{*}$. For $W\in\mathcal{U}$, $W$ is invertible and we have $W\in\mathcal{M}^{*}$ if and only if $\mathcal{A}^{*}(\mathcal{A}(WW^{\top})-y)=0$. The function $H(W)=\mathcal{A}^{*}(\mathcal{A}(WW^{\top})-y)$ is therefore a local defining function for $\mathcal{M}^{*}$. Its differential at $\overline{W}$ satisfies for $U\in\mathbb{R}^{n\times n}$,
$DH(\overline{W})[U]=\mathcal{A}^{*}A(\overline{W}U^{\top}+U\overline{W}^{\top}).$
@@
In the last two examples, the set $\nabla F^{-1}(0)$ has singular points, and so Assumption 2 does not hold globally. However, we showed that it holds on “most of the space”, as there exists a negligible set $\mathcal{S}$ such that $\mathcal{M}^{*}=\nabla F^{-1}(0)\setminus\mathcal{S}$ is a smooth manifold.
@@
Consider then the Riemannian gradient flow $w^{\circ}$ of the $\ell_{2}$ norm on $\mathcal{M}^{*}$ initialized at $w^{\mathrm{GF}}_{\infty}$. For any time horizon $T$ such that the trajectory of $w^{\circ}$ stays in $\mathcal{U}$ on the interval $[0,T]$, we can restrict our analysis to this local region, where our assumptions are satisfied. We can then invoke Proposition 2 to conclude that $\tilde{w}^{\lambda}$ converges to $w^{\circ}$ uniformly on intervals of the form $[\epsilon,T]$.
However, a key limitation arises when analyzing the long-time behavior: the results characterizing the limit points ( Proposition 4) do not apply if $w^{\circ}$ converges to a singular point outside $\mathcal{M}^{*}$. This situation can occur, as singular points might correspond to points that minimize the $\ell_{2}$ norm on $\mathcal{M}^{*}$ (e.g., sparse vectors for diagonal networks, or low-rank matrices for matrix sensing). Establishing convergence of $w^{\lambda}$ to such singular points remains an open and challenging problem, which we leave for future work.
@@ D.1 Additional Experimental Details
In all our figures and to align with the continuous-time analysis, training iterations refers to the rescaled "training time" $t_{k}=\gamma k$, where $k$ is the number of gradient steps and $\gamma$ the gradient descent stepsize. We run gradient descent $10^{7}$ iterations for Figure 2 and $10^{6}$ iterations for Figure 3.
@@
Experimental setup (Figure 4). We train a two-layer diagonal linear network of the form $f_{w}(x)=\langle u\odot v,\varphi(x)\rangle$, where $w=(u,v)\in\mathbb{R}^{2d}$ and $\odot$ denotes element-wise multiplication, on a 1D toy dataset. The input $x\in\mathbb{R}$ is mapped to a high-dimensional feature space via the feature map $\varphi(x)=\left[1,\cos\left(\tfrac{\pi x}{2}\right),\dots,\cos\left(\tfrac{\pi d_{f}x}{2}\right),\sin\left(\tfrac{\pi x}{2}\right),\dots,\sin\left(\tfrac{\pi d_{f}x}{2}\right)\right],$ with $d_{f}=30$. The teacher function is a sparse Fourier series $f(x)=1+\cos\left(\tfrac{6\pi x}{2}\right)+\sin\left(\tfrac{21\pi x}{2}\right)$ and is shown as a dotted light blue curve in Figure 4. The training dataset consists of $n=12$ input-output pairs $(x_{i},y_{i})$, where $x_{i}$ are sampled uniformly in $[-1,1]$ and $y_{i}=f(x_{i})$. These training points are shown as crosses in Figure 4. We optimise the squared loss $F(w)=\tfrac{1}{2n}\sum_{i=1}^{n}\left(y_{i}-f_{w}(x_{i})\right)^{2}$ using gradient descent with weight decay $\lambda=10^{-4}$. Finally, the initial weights are sampled from a centered Gaussian of variance $0.1$.
@@
Explaining the observed grokking phenomenon. At time $t_{1}=0$, the weights are randomly initialised and the training loss is high. By $t_{2}=10^{2}$, the training loss has dropped to nearly zero, and the iterates closely approximate the solution that would be obtained by unregularised gradient flow. This solution is fully characterised by the implicit regularisation result of [Woodworth et al., 2020], and it does not have a low norm. [^4] One could also reach the solution observed at time $t_{3}=10^{5}$ without using weight decay by employing a much smaller initialisation scale Woodworth et al. [2020], but at the cost of longer training time. Subsequently, around time $t=1/\lambda$, the weight norms begin to decrease, and by $t_{3}\approx 10^{5}$, they converge to the minimum-norm solution $(u^{\star},v^{\star})=\arg\min_{F(u,v)=0}\|u\|_{2}^{2}+\|v\|_{2}^{2}$. A straightforward calculation shows that the elementwise product $\beta^{\star}\coloneqq u^{\star}\odot v^{\star}$ solves the problem $\arg\min_{\langle\beta,x_{i}\rangle=y_{i}\forall i}\|\beta\|_{1}$. This is an $\ell_{1}$-minimisation problem, which (under RIP conditions) is known to recover the sparsest solution [Candes, 2008], explaining the zero test loss after the grokking phenomenon.

### v2

@@ B.4 Bounding the trajectories
If Assumptions˜1 and 2 hold, there exists a compact $K$ of $\mathbb{R}^{d}$ such that for any $\lambda>0$ and $t\geq 0$, $w^{\lambda}(t)\in K$. In particular, $\lim_{t\to\infty}w^{\lambda}(t)$ exists for any $\lambda>0$.
@@
Similarly to the proof of Lemma˜1, we can consider a neighbourhood $U$ of $\mathcal{M}$ where the PL inequality holds:
@@
1) For some fixed $\varepsilon>0$, Lemma˜2 then implies there is a $\lambda^{\star}>0$ and times $t(\lambda)$ such that for any $\lambda\in(0,\lambda^{\star})$ both hold
@@
Moreover, the proof of Lemma˜2 also implies that there is some compact $K_{1}$ of $\mathbb{R}^{d}$ such that for any $\lambda<\lambda^{\star}$ and $t\leq\frac{t(\lambda)}{\lambda}$, $w^{\lambda}(t)\in K_{1}$.
@@
Define $t^{\prime}=\min(\frac{t(\lambda)}{\lambda}+\frac{-\ln(\lambda)}{\eta},T_{\lambda})$. Using Equation˜18 for any $t\in(\frac{t(\lambda)}{\lambda},t^{\prime}]$:
@@
$\left\|w^{\lambda}(t)-w^{\lambda}(\frac{t(\lambda)}{\lambda})\right\|$ $\leq\int_{\frac{t(\lambda)}{\lambda}}^{t^{\prime}}\|\nabla F(w^{\lambda}(s))\|\mathrm{d}s+\lambda\int_{\frac{t(\lambda)}{\lambda}}^{t^{\prime}}\|w^{\lambda}(s)\|\mathrm{d}s$
$\leq\beta\int_{\frac{t(\lambda)}{\lambda}}^{t^{\prime}}\sqrt{F(w^{\lambda}(s))-F^{\star}}\mathrm{d}s-\frac{\lambda\ln(\lambda)}{\eta}(\|\Phi(w_{0})\|+\delta).$
From there, Equation˜20 yields for any $t\in(\frac{t(\lambda)}{\lambda},t^{\prime}]$:
$\left\|w^{\lambda}(t)-w^{\lambda}(\frac{t(\lambda)}{\lambda})\right\|$ $\leq\beta\int_{0}^{t^{\prime}-\frac{t(\lambda)}{\lambda}}\sqrt{\varepsilon}e^{-\frac{\eta}{2}s}\mathrm{d}s+\beta(t^{\prime}-\frac{t(\lambda)}{\lambda})\sqrt{\lambda(\frac{R_{1}}{\eta}+\frac{1}{2}R^{2})}-\frac{\lambda\ln(\lambda)}{\eta}(\|\Phi(w_{0})\|+\delta)$
$\leq\beta\sqrt{\varepsilon}\frac{2}{\eta}-C\sqrt{\lambda}\ln(\lambda)-\frac{\lambda\ln(\lambda)}{\eta}(\|\Phi(w_{0})\|+\delta),$
@@
3) Since $^{\prime}<T_{\lambda}, t=t$$^{\prime}=\frac{t(\lambda)}{\lambda}+\frac{-\ln(\lambda)}{\eta}$ by definition and
@@
Now define $\tilde{T}_{\lambda}=\inf\left\{t\geq\frac{t(\lambda)}{\lambda}\mid w^{\lambda}(t)\not\in U\right\}$. Since $\mathcal{M}$ minimizes $F$ on $U$, Equation˜21 implies by continuity that for any $t\in[t^{\prime},\tilde{T}_{\lambda}]$:
@@
In particular, for $K_{2}=B(0,\frac{2R_{1}}{\eta}+R^{2}+3\varepsilon)$ and any $t\in[t^{\prime},\tilde{T}_{\lambda}]$, $w^{\lambda}(t)\in K_{2}$. By continuity and compactness, $\inf_{w\in(\partial U)\cap K_{2}}F(w)>F^{\star}$ thanks to Lemma˜5. In consequence, we can choose $\lambda$ small enough so that Equation˜21 implies that for any $t\in[t^{\prime},\tilde{T}_{\lambda}]$,
@@
Assume now that $\tilde{T}_{\lambda}<\infty$. Since $w^{\lambda}(\tilde{T}_{\lambda})\in K_{2}$, the previous inequality implies by continuity that $w^{\lambda}(\tilde{T}_{\lambda})\not\in\partial U$, i.e., $w^{\lambda}(\tilde{T}_{\lambda})\in\mathring${U}. This however contradicts the definition of $\tilde{T}_{\lambda}$, so that $\tilde{T}_{\lambda}=\infty$. In particular for any $t\geq t^{\prime}$, $w^{\lambda}(t)\in K_{2}$.
To summarize, we have showed that there exists a small enough $\lambda^{\star}$, such that for any $\lambda\leq\lambda^{\star}$:
@@
As a consequence, the definability assumption of $F_{\lambda}$ along with the boundedness implies that $\lim_{t\to\infty}w^{\lambda}(t)$ exists thanks to kurdyka1998gradient.
B.5 Proof of Lemma˜3
@@
We consider the neighbourhood $U$ defined as in Lemma˜7. By definition, $F$ is constant on the manifold $\mathcal{M}$ and denote its value $F^{\star}$, i.e., $F^{\star}=\sup_{w\in\mathcal{M}}F(w)=\inf_{w\in\mathcal{M}}F(w)$.
@@
Thanks to Lemma˜5, for any $w\in U\setminus\mathcal{M}$, $F(w)>F^{\star}$. This implies the third condition,
@@
The arguments of Lemma˜7 extend to any family of solutions $(u^{\lambda})_{\lambda>0}$ satisfying the assumptions of Lemma˜3. In consequence, we can consider a compact $K$ of $\mathbb{R}^{d}$ such that for any $t\geq 0$ and $\lambda>0$, $u^{\lambda}(t)\in K$. From there, note again that $\inf_{w\in(\partial U)\cap K}F(w)>F^{\star}$.
@@
By monotonicity of the objective over time, $F(u^{\lambda}(t))+\lambda\|u^{\lambda}(t)\|^{2}_{2}<\min(F^{\star}+\varepsilon,\inf_{w\in(\partial U)\cap K}F(w))$ for any $t\geq 0$. Since $u^{\lambda}(t)$ is continuous, it implies that $u^{\lambda}(t)\in\overset{\circ}{U_{\varepsilon}}$ for any $t\geq 0$, which concludes the proof of Lemma˜3. ∎
@@ B.6 Strict Minimality
Under the same assumptions than Proposition˜4 with $w^{\star}=\lim_{t\to\infty}\tilde{w}^{\circ}(t)$, there exists a $\delta^{\star}>0$ such that for any $\delta\in(0,\delta^{\star})$, there exists $\varepsilon>0$ and $\lambda^{\star}>0$ such that for any $\lambda\in(0,\lambda^{\star})$
@@
Let $\delta^{\star}>0$ be such that for any $\in\mathcal{M}\cap B(w^{\star},2\delta^{\star}), ww$$\neq w^{\star}\implies\|w\|^{2}>\|w^{\star}\|^{2}$ and such that Equation˜14 holds. Now let $\delta\in(0,\delta^{\star})$. We now fix $\lambda^{\star}>0$ arbitrarily small, choose $\lambda\in(0,\lambda^{\star})$ and define
@@ Appendix C Applications
In this section, we provide additional details to the examples discussed in Section˜5, and specify how our theoretical results can be applied in various settings.
@@
Denote the solution of minimal $\ell_{2}$ norm with $w^{\star}=X^{+}y$, where $X^{+}$ is the Moore-Penrose pseudoinverse of $X$. The problem is convex and the critical set of $F$ is the affine subspace $\mathcal{M}=w^{\star}+\mathrm{Ker}(X)$, which is a manifold: Assumption˜2 is satisfied.
@@
DLNs serve as a toy example to understand the influence of the architecture on the training dynamics of neural networks [pesme2024deep]. The corresponding optimization problem writes
@@
$\begin{split}u\odot\left[X^{\top}(X(u^{2}-v^{2})-y)\right]&=0,\\ v\odot\left[X^{\top}(X(u^{2}-v^{2})-y)\right]&=0\end{split}$ (25)
@@
The set $\mathcal{M}^{*}=\nabla F^{-1}(0)\cap(\mathbb{R}^{*})^{2d}$ is a smooth manifold of dimension 2dn2d-n.
@@
Let $(\bar{u},\bar{v})\in\mathcal{M}^{*}$. Denote $W$ a neighborhood of $(\bar{u},\bar{v})$ such that $U\subset(\mathbb{R}^{*})^{2d}$. The function $H:\mathbb{R}^{2d}\rightarrow R^{d}$ with $H(u,v)=X^{\top}(X(u^{2}-v^{2})-y)$ is a local defining function for $\mathcal{M}^{*}$, in the sense that for $(u,v)\in W$, we have $(u,v)\in\mathcal{M}^{*}\iff H(u,v)=0$.
@@
It is clear that, since all coordinates of $(\bar{u},\bar{v})$ are nonzero, the map $(\Delta u,\Delta v)\mapsto\bar{u}\odot\Delta u-\bar{v}\odot\Delta v$ is a surjection on $\mathbb{R}^{d}$, and therefore $\text{rank}(DH(\bar{u},\bar{v}))=\text{rank}(X^{\top}X)=n$. This proves that $\mathcal{M}^{*}$ is a manifold of dimension 2dn2d-n [boumal2023introduction, §3.2]. ∎
Because of the singular points in $\mathcal{M}$, the function $F$ does not satisfy Assumption˜2 globally. However, our results can still be applied locally: see the paragraph below for details.
@@
A typical example is symmetric matrix completion, where the goal is to recover an unknown matrix $M^{*}\in\mathbb{R}^{n\times n}$ from a subset of observed entries with coefficients in $\Omega\in\{1\dots n\}^{2}$: the objective function writes $F(W)=\sum_{(i,j)\in\Omega}\left((WW^{\top})_{ij}-M_{ij}^{*}\right)^{2}$. Note that the asymmetric case presented in Section˜5, Equation˜8, can also be written as a symmetric matrix completion problem, by setting
@@
Let $\overline{W}\in\mathcal{M}^{*}$, and let $\mathcal{U}$ a neighborhood of $\overline{W}$ such that $\mathcal{U}\subset\mathbb{R}^{n\times n}_{*}$. For $\in\mathcal{U}, WW$ is invertible and we have $W\in\mathcal{M}^{*}$ if and only if $\mathcal{A}^{*}(\mathcal{A}(WW^{\top})-y)=0$. The function $H(W)=\mathcal{A}^{*}(\mathcal{A}(WW^{\top})-y)$ is therefore a local defining function for $\mathcal{M}^{*}$. Its differential at $\overline{W}$ satisfies for $U\in\mathbb{R}^{n\times n}$,
$DH(\overline{W})[U]=\mathcal{A}^{*}\mathcal{A}(\overline{W}U^{\top}+U\overline{W}^{\top}).$
@@
In the last two examples, the set $\nabla F^{-1}(0)$ has singular points, and so Assumption˜2 does not hold globally. However, we showed that it holds on “most of the space”, as there exists a negligible set $\mathcal{S}$ such that $\mathcal{M}^{*}=\nabla F^{-1}(0)\setminus\mathcal{S}$ is a smooth manifold.
@@
Consider then the Riemannian gradient flow $w^{\circ}$ of the $\ell_{2}$ norm on $\mathcal{M}^{*}$ initialized at $w^{\mathrm{GF}}_{\infty}$. For any time horizon $T$ such that the trajectory of $w^{\circ}$ stays in $\mathcal{U}$ on the interval $[0,T]$, we can restrict our analysis to this local region, where our assumptions are satisfied. We can then invoke Proposition˜2 to conclude that $\tilde{w}^{\lambda}$ converges to $w^{\circ}$ uniformly on intervals of the form $\epsilon,T]$.
However, a key limitation arises when analyzing the long-time behavior: the results characterizing the limit points ( Proposition˜4) do not apply if $w^{\circ}$ converges to a singular point outside $\mathcal{M}^{*}$. This situation can occur, as singular points might correspond to points that minimize the $\ell_{2}$ norm on $\mathcal{M}^{*}$ (e.g., sparse vectors for diagonal networks, or low-rank matrices for matrix sensing). Establishing convergence of $w^{\lambda}$ to such singular points remains an open and challenging problem, which we leave for future work.
@@ D.1 Additional Experimental Details
In all our figures and to align with the continuous-time analysis, training iterations refers to the rescaled "training time" $t_{k}=\gamma k$, where $k$ is the number of gradient steps and $\gamma$ the gradient descent stepsize. We run gradient descent $10^{7}$ iterations for Figure˜2 and $10^{6}$ iterations for Figure˜3.
@@
Experimental setup (Figure˜4). We train a two-layer diagonal linear network of the form $f_{w}(x)=\langle u\odot v,\varphi(x)\rangle$, where $w=(u,v)\in\mathbb{R}^{2d}$ and $\odot$ denotes element-wise multiplication, on a 1D toy dataset. The input $x\in\mathbb{R}$ is mapped to a high-dimensional feature space via the feature map $\varphi(x)=\left[1,\cos\left(\tfrac{\pi x}{2}\right),\dots,\cos\left(\tfrac{\pi d_{f}x}{2}\right),\sin\left(\tfrac{\pi x}{2}\right),\dots,\sin\left(\tfrac{\pi d_{f}x}{2}\right)\right],$ with $d_{f}=30$. The teacher function is a sparse Fourier series $f(x)=1+\cos\left(\tfrac{6\pi x}{2}\right)+\sin\left(\tfrac{21\pi x}{2}\right)$ and is shown as a dotted light blue curve in Figure˜4. The training dataset consists of $n=12$ input-output pairs $(x_{i},y_{i})$, where $x_{i}$ are sampled uniformly in [1,1][-1,1] and $y_{i}=f(x_{i})$. These training points are shown as crosses in Figure˜4. We optimise the squared loss $F(w)=\tfrac{1}{2n}\sum_{i=1}^{n}\left(y_{i}-f_{w}(x_{i})\right)^{2}$ using gradient descent with weight decay $\lambda=10^{-4}$. Finally, the initial weights are sampled from a centered Gaussian of variance $0.1$.
@@
Explaining the observed grokking phenomenon. At time $t_{1}=0$, the weights are randomly initialised and the training loss is high. By $t_{2}=10^{2}$, the training loss has dropped to nearly zero, and the iterates closely approximate the solution that would be obtained by unregularised gradient flow. This solution is fully characterised by the implicit regularisation result of [woodworth2020kernel], and it does not have a low norm. [^5] One could also reach the solution observed at time $t_{3}=10^{5}$ without using weight decay by employing a much smaller initialisation scale [woodworth2020kernel], but at the cost of longer training time. Subsequently, around time $t=1/\lambda$, the weight norms begin to decrease, and by $t_{3}\approx 10^{5}$, they converge to the minimum-norm solution $(u^{\star},v^{\star})=\arg\min_{F(u,v)=0}\|u\|_{2}^{2}+\|v\|_{2}^{2}$. A straightforward calculation shows that the elementwise product $\beta^{\star}\coloneqq u^{\star}\odot v^{\star}$ solves the problem $\arg\min_{\langle\beta,x_{i}\rangle=y_{i}\forall i}\|\beta\|_{1}$. This is an $\ell_{1}$-minimisation problem, which (under RIP conditions) is known to recover the sparsest solution [candes2008restricted], explaining the zero test loss after the grokking phenomenon.
Appendix E Heuristic analysis on how small $\lambda$ needs to be for grokking to emerge
Note that our theoretical results are derived in the asymptotic regime $\lambda\to 0$, since this setting allows for a tractable and general analysis. Extending the theory to obtain explicit results for a fixed $\lambda>0$ is considerably more challenging. That said, we can offer some intuition regarding how small $\lambda$ needs to be for grokking to emerge.
Grokking depends on a clear separation between two phases: an initial phase where the iterates converge and stagnate at the solution of the unregularised gradient flow, and a second phase driven by weight decay, during which test performance improves. For grokking to be observable, the regularised gradient flow should approach the unregularised limit before weight decay begins to significantly influence the dynamics.
To formalize this intuition, we can define two characteristic times: $t_{\rm GF}$, the convergence time of the unregularized gradient flow, measured as the time at which the gradient norm substantially decreases relative to its initial value; and a second time $t_{\rm WD}$ of order $1/\lambda$, associated with the onset of the regularization effects. When $\lambda$ is small enough that $t_{\rm GF}\ll 1/\lambda$, we expect to observe grokking-like behavior. Specifically, for some threshold $\varepsilon\ll 1$ (e.g., $\varepsilon=0.01$): let $t_{\rm GF}$ such that $\|\nabla F(w_{t_{\rm GF}})\|\approx\varepsilon\|\nabla F(w_{0})\|$. Now let $t_{\rm WD}$ denote the time when weight decay kicks in: i.e. when the magnitude of the unregularised gradient becomes comparable to the magnitude of the weight decay term: $\|\nabla F(w_{t_{\rm WD}})\|\approx\lambda\|w_{t_{\rm WD}}\|$. Since at time $t_{\rm WD}$, the solution is close to the gradient flow solution $w^{\rm GF}$, we can consider $\|w_{t_{\rm WD}}\|\approx\|w^{\rm GF}\|$. The condition for grokking to occur (i.e., a plateau in test loss followed by an improvement of the test loss) is thus that $t_{\rm GF}\ll t_{\rm WD}$. Translating this condition in terms of gradients, we obtain: $\|\nabla F(w_{t_{\rm GF}})\|\gg\|\nabla F(w_{t_{\rm WD}})\|$, which, using the approximations above, implies: $\varepsilon\|\nabla F(w_{0})\|\gg\lambda\|w^{GF}\|$. Simplifying further (absorbing $\varepsilon$ into a constant), we have the practical guideline: $\lambda\ll\frac{\|\nabla F(w_{0})\|}{\|w^{\rm GF}\|}$. Hence, grokking occurs when the weight decay parameter $\lambda$ is sufficiently small compared to the ratio between the initial gradient magnitude and the norm of the unregularised gradient flow solution.